OpenStudy (anonymous):

The ratio of two side lengths of the triangle is given. Solve for the variable, then find the length of AB. AB:CB is 3:4 please help !

OpenStudy (anonymous):



OpenStudy (e.mccormick):

They give you a ratio and two equations. Do you know how you would use the ratio to set up the equations equal to each other?

OpenStudy (anonymous):

x/(x-2)=3/4 simplify and find x

OpenStudy (anonymous):

i honestly dont understand what to do.... :/

OpenStudy (anonymous):

AB:BC =AB/BC plug in values and simplify

OpenStudy (e.mccormick):

It is like saying \(\cfrac{1}{2}=\cfrac{2}{4}\). Only now you have equations in the mix. That means you set up the equality like ArslanQureshi did, then simplify it.

OpenStudy (anonymous):

how do you set it up cuz i think tht i can do the rest if i start from there mainly cuz idk how to set stuff up but im alright with the actual math part.

OpenStudy (anonymous):

okay AB:BC means AB/BC in its simplest form see e.mccormick's post above if AB is 2 and BC is 4 then AB:BC becomes 1:2 which equals 1/2. So in questions like these write ratio as A/B and put it equal to given expressions of A and B e.g. IF A:B = 1:3 AND A=X B=2X THEN 1/3=X/2X WHICH SIMPLIFYS TO X=1 get it??

OpenStudy (anonymous):

wow yes lol thankyou!

OpenStudy (e.mccormick):

Which is where his setup comes from:\[\frac{x}{(x-2)}=\frac{3}{4}\]And why solving that will answer your question. So start by cross multiplying and go from there.